# MPD-SGR: Robust Spiking Neural Networks with Membrane Potential Distribution-Driven Surrogate Gradient Regularization

### v1

@@ Methods
We further evaluate the impact of MPD-SGR on the robustness of SNNs under non-gradient attacks, specifically two types of random perturbations: Gaussian ($\delta\sim\mathcal{N}(0,\epsilon)$) and Uniform ($\delta\sim\mathcal{U}(-\epsilon,\epsilon)$) noise. All results are reported using a consistent noise intensity of $\epsilon=0.1$. As shown in Table 3, our method maintains high classification performance under both perturbation types, significantly outperforming the REG baseline, which suffers severe performance degradation (accuracy drops of 48.09% and 23.86%). Although SR exhibits a degree of robustness, its substantially lower clean accuracy (66.76%) fundamentally constrains its performance under noisy conditions, resulting in absolute accuracies below those achieved by our approach. These results indicate that MPD-SGR exhibits strong robustness to non-gradient perturbations, showcasing its versatility across diverse input disturbances.
@@
We investigate the applicability of our method to other commonly used SG functions, including Rectangular (Rec) (wu2019direct), Sigmoid (Sig) (roy2019scaling) and Superspike (Sup) (zenke2018superspike). As presented in Table 4, our proposed MPD-SGR method consistently enhances robustness across different SG functions, further confirming the versatility and effectiveness of our methods.
@@
Considering the significant impact of input spike encoding method on the robustness of SNNs, we also compare the performance of the MPD-SGR method with the vanilla SNN model under different spike coding methods, including Direct Coding (DIR), Poisson Coding (POS) and Randomized Smoothing Coding (RSC) (wu2024rsc). The proposed MPD-SGR method exhibits consistent performance superiority over the baseline SNN model across all tested spike coding schemes (Table 5), demonstrating the compatibility of our method with various spike coding approaches.
@@ Conclusion
This paper presents a theoretical and empirical investigation into the role of neuronal membrane potential distribution (MPD) and surrogate gradient (SG) mechanisms in enhancing adversarial robustness of deep spiking neural networks (SNNs). Through rigorous gradient analysis, we establish a novel theoretical framework that formally connects input perturbation sensitivity with the magnitude characteristics of SG in SNNs. Our analysis reveals that reducing the proportion of membrane potentials within the gradient-available range of the SG function effectively mitigates this input perturbation sensitivity of SNNs.
Based on these theoretical insights, we propose MPD-driven surrogate gradient regularization (MPD-SGR), which systematically constrains the MPD during training to optimize the gradient flow, thereby enhancing the network’s resilience against adversarial attacks. Extensive experiments across multiple benchmark datasets and SNN architectures show that MPD-SGR maintains comparable clean accuracy while consistently improving robustness against gradient-based attacks and random noise, outperforming existing defense methods. The proposed approach also demonstrates remarkable adaptability across diverse network configurations, SG function variants, and spike coding schemes. These results position MPD-SGR as a promising, generalizable defense strategy for SNNs under diverse perturbation scenarios.
@@ Acknowledgements
This work was supported by the National Natural Science Foundation of China (Grant Nos. 62236007, 62276235, and 32441113) and the National Key Research and Development Program of China (Grant No. 2024YDLN0005). The authors would also like to acknowledge anonymous reviewers and chairs for providing insightful comments to help improve this work.

### v2

@@ Methods
We further evaluate the impact of MPD-SGR on the robustness of SNNs under non-gradient attacks, specifically two types of random perturbations: Gaussian ($\delta\sim\mathcal{N}(0,\epsilon)$) and Uniform ($\delta\sim\mathcal{U}(-\epsilon,\epsilon)$) noise. All results are reported using a fixed noise intensity of $\epsilon=0.1$. As shown in Table 3, our method maintains high classification performance under both perturbation types, significantly outperforming the REG baseline, which suffers severe performance degradation (accuracy drops of 48.09% and 23.86%). Although SR exhibits a degree of robustness, its substantially lower clean accuracy (66.76%) fundamentally constrains its performance under noisy conditions, resulting in absolute accuracies below those achieved by our approach. These results indicate that MPD-SGR exhibits strong robustness to non-gradient perturbations, showcasing its versatility across diverse input disturbances.
@@
We investigate the applicability of our method to other commonly used SG functions, including Rectangular (Rec) (wu2019direct), Sigmoid (Sig) (roy2019scaling) and Superspike (Sup) (zenke2018superspike). As presented in Table 4, our proposed MPD-SGR method consistently enhances robustness across different SG functions, further confirming the versatility and effectiveness of our method.
@@
Considering the significant impact of input spike encoding methods on the robustness of SNNs, we also compare the performance of MPD-SGR with that of the vanilla SNN model under different spike coding methods, including Direct Coding (DIR), Poisson Coding (POS) and Randomized Smoothing Coding (RSC) (wu2024rsc). The proposed MPD-SGR method exhibits consistently superior performance over the baseline SNN model across all tested spike coding schemes (Table 5), demonstrating the compatibility of our method with various spike coding approaches.
@@ Conclusion
This paper presents a theoretical and empirical investigation into the role of neuronal membrane potential distribution (MPD) and surrogate gradient (SG) mechanisms in enhancing the adversarial robustness of deep spiking neural networks (SNNs). Through rigorous gradient analysis, we establish a theoretical framework that formally connects input perturbation sensitivity to the magnitude characteristics of SG in SNNs. Our analysis reveals that reducing the proportion of membrane potentials within the gradient-available range of the SG function effectively mitigates the sensitivity of SNNs to input perturbations. Based on these theoretical insights, we propose MPD-driven surrogate gradient regularization (MPD-SGR), which systematically constrains the MPD during training to optimize gradient flow, thereby enhancing the network’s resilience against adversarial attacks. Extensive experiments across multiple benchmark datasets and SNN architectures show that MPD-SGR maintains comparable clean accuracy while consistently improving robustness against gradient-based attacks and random noise, outperforming existing defense methods. The proposed approach also demonstrates remarkable adaptability across diverse network configurations, SG function variants, and spike coding schemes. These results position MPD-SGR as a promising, generalizable defense strategy for SNNs under diverse perturbation scenarios.
@@ Acknowledgements
This work was supported by the National Natural Science Foundation of China (Grant Nos. 62236007, 32441113, and 62276235). The authors would also like to acknowledge anonymous reviewers and chairs for providing insightful comments to help improve this work.